# Magneto-optical trap loading with an effusive oven in    a large optical access experiment

### v1

@@ B. Zeeman slower
The ZS beam enters through a heated sapphire window (at $\,{}^{\circ}$, to prevent deposition of strontium), as seen at the right of Fig. 1(a). It has a power of 70 mW (taking into account a non-negligible loss through the window), a $1/e^{2}$ radius of $4$ mm, and a detuning of $-370$ MHz. The slower design is based on using permanent magnets that are placed inside of cups surrounding opposite sides of the nipple connecting the 2D OM and 2D MOT chambers (50 cm apart) [44]. It produces a transverse magnetic field (xy plane) that increases non-monotonically from $-400$ G to 00 G over the total length of 27 cm. We note that an iron plate attached at the slower exit provides a magnetic shielding of the sensitive end field (mainly against the 2D MOT field).
For our simulations, we have derived a heuristic model (Eq. A8), where the ZS magnetic field is pieced together by linear segments with given start and end point values. Note that due to its transverse nature, the inclusion of light polarization is essential for a correct description of the slowing. For a smooth test profile that closely matches the experimental one, we choose $n_{S}=30$ segments in our
simulations. As for the 2D OM (Sec. II.A), the effect of the ZS is evaluated at the 2D MOT, described next, while Sec. IV considers the numerical optimization.
@@
To realize the 2D MOT, we use permanent magnets creating a $67$ G/cm field gradient along the axes of two retroreflected beams with an individual power of 20 mW, a $1/e^{2}$ radius of $3.2$ mm, and a detuning of $-32$ MHz. The magnets are placed inside of holders [not shown in Fig. 1(a)] above and below (z axis) the jet entrance and exit nipples of the 2D MOT chamber, with the magnetization directions being opposite (y axis) on the respective nipples. The resulting magnetic field is zero along the axis perpendicular to the main 2D MOT plane (xy plane) as well as non-cylindrically-symmetric around this axis (unlike in a 3D MOT). The atoms are transferred into the 3D MOT along this axis (Sec. II.D).
The permanent magnet field of the 2D MOT is modeled accordingly by Eq. A7, where we include the experimental distances between the magnets (modeled as dipoles), their volumes ($V=1^{\prime\prime}\times 1/2^{\prime\prime}\times 1/4^{\prime\prime}$ each) and remanent magnetization ($B_{0}=1.44$ T; N52 grade neodymium magnet). Due to the particular asymmetry of the field in the main 2D MOT plane, the inclusion of light polarization is essential for a correct description of the 2D MOT loading. We report in Tab. 1 the loading rates obtained experimentally and numerically to evaluate the effects of the 2D OM and the ZS magnets.
2D MOT loading rates (atoms/s ×$\times 10^{8}$)
@@
-- ZS, -- 2D OM $1.6$ $1.3$ $19\,\%$
-- ZS, $+$ 2D OM $3.6$ $2.1$ $42\,\%$
$+$ ZS, -- 2D OM $7.1$ $9$ $-27\,\%$
$+$ ZS, $+$ 2D OM $15$ $13$ $13\,\%$
Table 1: A comparison of the 2D MOT loading rates obtained experimentally ($\mathcal{F}_{exp}$) and numerically ($\mathcal{F}_{sim}$). ”$+$” and ”--” refer to the 2D OM or the ZS magnets being used and not used, respectively.
Starting with row/case 1, where the atom dynamics are not affected by the 2D OM or the ZS magnets (note that the ZS beam itself is essential for the loading), the discrepancy between the experiment and simulation is less than $20\,\%$. This is noteworthy considering there is only one free numerical parameter (the experimental flowrate).
The simulation is highly influenced by the initial velo-
city profile, and by employing a geometrical reasoning (Sec. II.A) proves to yield closely matching results. The atom lifetime is also an important factor to consider (App. A2), given that the atoms can get shelved to a metastable state as they travel through the Zeeman nipple and hence elude the 2D MOT capture.
@@
[…]
a detuning of $-30$ MHz. They are lowered by approximately one beam radius with respect to the quadrupole field zero, resulting in an increased loading efficiency (discussed in detail in Sec. III). Additionally, by mirroring them at $25\,^{\circ}$ with respect to the radial axis perpendicular to the objectives axis (i.e., with respect to the y axis) grants the necessary optical access for the high NA optical tweezers while simultaneously covering most of the science region for an enhanced loading (see Sec. IV for the numerical verification). One of the retroreflected pairs also contains the repumping light with a total power of $8$ mW; we find the repumping to be critical as most of the atoms get shelved to the metastable state $5s5p\,^{3}\!P_{2}$ before reaching the science region.
The 3D MOT magnetic field is modeled accordingly by Eq. A6, and the radiation pressure forces corresponding to the radial plane beams take into account them being lowered (1 radial beam radius) as well as the angling ($25\,^{\circ}$). Note that, for completeness, this field retains a full 3D character, such that its effect is present in the transfer region (the same is true for the 2D MOT field). This information feeds into the scattering cross-sections (Eq. A11), thus affecting the atom dynamics. The repumping is present in one of the radial beam pairs, like in the experiment. As mentioned in the introduction of Sec. II, when a given superparticle drifts into the repumping region, it is revived with a new lifetime that is enhanced compared to the non-repumping region.
@@ III. Main results
Our 3D MOT loading involves a simple technique based on moving molasses: With the axial 3D MOT beams intersecting the 2D MOT, a cold flux loads the 3D MOT by introducing a radiation-pressure imbalance between them.
This imbalance is due to a difference in detunings, which for a broad range of parameters (discussed below) does not cause the 3D MOT to be emptied. The atoms instead accumulate below the quadrupole field zero as set by the degree of the imbalance. With this technique, an advantageous MOT laser beam configuration is obtained as compared to adding a push/cold flux beam, which may otherwise require operating the MOT beams at angles compromising the loading [22, 23]. Moreover, a broad optical access to the scientific study region is achieved, creating additional laser-pathways for atom array manipulation with a high NA optical tweezer setup.
Below in this section, we  first  develop  an  analytical
model for the 3D MOT loading using our technique and then compare the main experimental observations to this model as well as the numerical model. Lastly, we discuss a configuration with a misaligned bottom beam, which creates the necessary radiation-pressure imbalance for loading with equally detuned axial beams. We note that this special configuration obfuscates analytical treatment and is thus not central to our present analysis.
@@ A. The analytical model
To gain an in-depth physical insight into our loading scheme (with overlapped axial beams), we develop an analytical model incorporating the atomic diffusion effect and the shelving to the metastable state $5s5p\,^{3}\!P_{2}$ during the transfer, as well as the loss to the dark state $5s5p\,^{3}\!P_{0}$ in the final trap region where repumping is present. The metastable-state shelving freezes the velocity distribution resulting from the diffusion, while the dark-state loss limits the loading. Assuming the atoms are transferred from the 2D MOT location with the flowrate $\mathcal{F}_{0}$, we obtain the following equation for the atom number in the 3D MOT:
$N=\mathcal{F}_{0}\times\left(\mathcal{P}_{\oplus}\times\mathcal{P}_{\ominus}\tau_{\ominus}\right)\quad\quad\quad\quad\;\,(1)$
where $\mathcal{P}_{\oplus}$ is the transfer probability including the diffusion effect and the metastable-state shelving; $\mathcal{P}_{\ominus}$ is the transfer probability including losses due to various MOT cut-offs (discussed below), and due to the confinement weakening as a result of a shift in the equilibrium position with respect to the radial beams center (caused by an axial beam detuning imbalance); and $\tau_{\ominus}$ is the time until the dark-state decay occurs in the trap.
For the transfer probabilities, we write
@@
$\mathcal{P}_{\oplus}$ $=H(d_{\textsf{c}}-d_{0})\times er\!f\left(\frac{d_{0}}{\sqrt{2\sigma_{r}^{2}}}\right)$
[…]
and
[…]
Figure 2: Color online. Diagrams showing the 3D MOT atom numbers (normalized; see text for absolute) using our loading technique connecting the 2D and 3D MOTs via moving molasses, obtained analytically, experimentally, and numerically. The experimental and numerical data points are obtained at the values seen on the axes, with the contours enhancing the features. The vertical dashed line marks the fixed bottom beam detuning ($-46$ MHz), and the slanted dashed line delineates approximately the sloped analytical feature. The cross and the dotted circle highlight respectively the location with the greatest analytical atom number ($44.8$ G/cm gradient at $-7.3$ MHz top beam detuning) and the greatest experimental as well as numerical atom number ($48$ G/cm gradient at $-10$ MHz top beam detuning).
The physical intuition behind these two equations is explained in what follows.
Firstly, Eq. 1a is seen to be a symmetric sum of products between the Heaviside functions $H$ and the error functions $er\!f$. The error functions are the probabilities for an atom to be confined by the 3D MOT after radially accumulating a diffusion-based Gaussian-position-spread
($\sigma_{r}^{2}$ variance) during the transfer from the 2D MOT to the 3D MOT. The Heaviside functions assigned to these error functions separate two regimes: First, where the field gradient is shallow enough for the radial confinement distance $d_{\textsf{c}}$ to extend beyond the distance $d_{0}$ that the radial beams travel inside the cell to its center (3D MOT), and, second, where $d_{\textsf{c}}$ is within $d_{0}$. (For $d_{\textsf{c}}=d_{0}$, $H(0)=1/2$ and continuity is preserved.) The unknown variables in Eq. 1a are as follows:
@@
(i) $d_{\textsf{c}}=\frac{3|\delta_{0}|}{\mu B^{\prime}}$, where $\frac{\delta_{0}}{2\pi}$ is the radial beam detuning, $\mu$ is the gyromagnetic ratio, and $B^{\prime}>0$ is the axial gradient (twice stronger than the radial). This expression is obtained by assuming for simplicity that the furthest capture occurs midway between the peak capture distance $\frac{2|\delta_{0}|}{\mu B^{\prime}}$ and twice this distance (where the radiation pressure force is significantly reduced). We note that, in general, the confinement distance has a complex dependence on different trapping parameters [58].
@@
(ii) $\sigma_{r}=H(\tau-t)\sigma_{r,t}+H(t-\tau)\sigma_{r,s}$, where the Heaviside functions separate the cases where the time $\tau$ until the metastable-state shelving is longer than the transfer time $t$ and vice versa; and $\sigma^{2}_{r,t}$ and $\sigma^{2}_{r,s}$ are the variances of the Gaussians describing the radial position spreads for these respective cases. When the shelving occurs, the corresponding speed Gaussian is assumed to freeze at $\tau$ and evolve the position Gaussian until $t$ by being in quadrature with the position Gaussian at $\tau$ and through a correlation term; thus, $\sigma_{r,s}=\sqrt{\sigma^{2}_{r,\tau}+\sigma^{2}_{v,\tau}(t-\tau)^{2}+2\rho\sigma_{r,\tau}\sigma_{v,\tau}(t-\tau)}$, where $\sigma^{2}_{r,\tau}$ and $\sigma^{2}_{v,\tau}$ are the variances of respectively the position and speed Gaussians at $\tau$, and $\rho=\frac{\sqrt{3}}{2}$ is the correlation between them (Brownian in nature). The remaining quantities seen in $\sigma_{r}$ are expressed as follows. The transfer time $t=\frac{d-d_{\textsf{e}}}{-v_{max}}>0$, where $d$ is the distance between the 2D and 3D MOTs; $d_{\textsf{e}}=\frac{\delta_{\text{b}}-\delta_{\text{t}}}{2\mu B^{\prime}}$ is the axial equilibrium position of the 3D MOT, as we assume the bottom and top beams of respective detunings $\frac{\delta_{\text{b}}}{2\pi}$ and $\frac{\delta_{\text{t}}}{2\pi}$ have the same intensity; and $v_{max}=\frac{\delta_{\text{b}}-\delta_{\text{t}}}{2k}$ is the maximal speed achieved during the transfer, assuming there is no magnetic field in the transfer region ($k=\frac{2\pi}{\lambda}$ is the angular wavenumber of the blue transition). The time until shelving, $\tau$, is obtained using Eq. A17 with $\mathcal{E}=1$ (no repumping) and $\Gamma_{23}^{\prime}$ as in Eq. A18, but we use here a 2-level model total saturation parameter for the bottom and top beams, $s_{\text{bt}}=\frac{I_{\text{bt}}/I_{sat}}{1+4\frac{(\delta_{\text{b}}-kv_{max})^{2}}{\Gamma^{2}}}+\frac{I_{\text{bt}}/I_{sat}}{1+4\frac{(\delta_{\text{t}}+kv_{max})^{2}}{\Gamma^{2}}}$ ($I_{\text{bt}}$
is the intensity of an axial beam). The variance $\sigma^{2}_{r,\tau}=\frac{2D_{r}\tau^{3}}{3^{2}m^{2}}$ ($\sigma^{2}_{r,t}$ uses $\tau\rightarrow t$), and $\sigma^{2}_{v,\tau}=\frac{2D_{r}\tau}{3m^{2}}$, where the transfer-region diffusion coefficient $D_{r}$ is obtained using Eq. A13 with the total saturation parameter being $s_{\text{bt}}$.
@@
Next, in Eq. 1b, we have the probability $\mathcal{P}_{\ominus}$ written in terms of three Heaviside functions and a Gaussian function. The leftmost Heaviside function is there to ensure that the MOT is empty when $\delta_{\text{t}}$ is positive and larger than $|\delta_{\text{b}}|$, as the sign of the confinement is then flipped. The other Heaviside functions ensure the MOT being empty when the radial beams do not encompass the MOT equilibrium position $d_{\textsf{e}}$ (as set by the axial beam imbalance), given the locations $d_{u}$ and $d_{l}$ of the upper and lower radial beam edges with respect to the quadrupole field zero. Note that similar Heaviside functions can be introduced to ensure an empty MOT when the transfer speed overcomes the axial capture speed and when $d_{\textsf{e}}$ exceeds the axial confinement distance. We omit these here as for our parameters the leftmost Heaviside function provides a stronger cut-off. Regarding the Gaussian in Eq. 1b, it represents an adjustment to the probability given a trapped atom at the position $d_{p}$ with respect to the center of the radial beams ($w_{0}$ waist radius) is not at their intensity maximum, where we assume the confinement to be the strongest. The unknown variables in Eq. 1b are as follows:
@@
The last quantity in Eq. 1, the time $\tau_{\ominus}$ until the dark-state decay occurs in the trap, is obtained using Eq. A17
with $\mathcal{E}=27$ (repumping present) and $\Gamma_{23}^{\prime}$ as in Eq. A18, but we use here a 2-level model total saturation parameter for all the 3D MOT beams, $s_{\text{all}}=s_{\text{bt,B}}+4s_{0}$, where $s_{\text{bt,B}}$ is the total saturation parameter given by $s_{\text{bt}}$ with the substitution $kv_{max}\rightarrow\mu B^{\prime}d_{\textsf{e}}$ (assuming a motionless atom), and $s_{0}=\frac{\frac{I_{0}}{I_{sat}}exp\left(-2\frac{d_{p}^{2}}{w_{0}^{2}}\right)}{1+4\frac{\delta_{0}^{2}}{\Gamma^{2}}}$ is the radial beam saturation parameter (assuming axial confinement) with $I_{0}$ being the peak radial beam intensity. We note that the introduction of $\tau_{\ominus}$ impacts greatly the physics of our observations, as discussed below.

### v2

@@ B. Zeeman slower
[…]
For our simulations, we have derived a heuristic model (Eq. A8), where the ZS magnetic field is pieced together by linear segments with given start and end point values. Note that due to its transverse nature, the inclusion of light polarization is essential for a correct description of the slowing. For a smooth profile that closely matches the experimental one, we choose $n_{S}=30$ segments in our simulations. In the Supplemental Material [49], we outline a method for the numerical optimization, which we note has not been implemented in the present case. Nonetheless, we expect it to yield accurate optimization results, given that the heuristic model for the ZS magnetic field leads to comparable numerical and experimental 2D MOT loading rates discussed below.
@@
To realize the 2D MOT, we use permanent magnets creating a $67$ G/cm field gradient along the axes of two retroreflected beams with an individual power of 20 mW, a $1/e^{2}$ radius of $3.2$ mm, and a detuning of 32-32 MHz. The magnets are placed inside of holders [not shown in Fig. 2(a)] above and below (z axis) the atomic beam entrance and exit nipples of the 2D MOT chamber, with the magnetization directions being opposite (y axis) on the respective nipples. The resulting magnetic field is zero along the axis perpendicular to the main 2D MOT plane (xy plane) as well as non-cylindrically-symmetric around this axis (unlike in a 3D MOT). The atoms are transferred into the 3D MOT along this axis (Sec. II.D).
The permanent magnet field of the 2D MOT is modeled accordingly by Eq. A7, where we include the experimental distances between the magnets (modeled as dipoles), their volumes ($V=1^{\prime\prime}\times 1/2^{\prime\prime}\times 1/4^{\prime\prime}$ each) and remanent magnetization ($B_{0}=1.44$ T; N52 grade neodymium magnet). Due to  the  particular  asymmetry
of the field in the main 2D MOT plane, the inclusion of light polarization is essential for a correct description of the 2D MOT loading. We report in Tab. 1 the loading rates obtained experimentally and numerically to evaluate the effects of the 2D OM and the ZS magnets.
2D MOT loading rates (atoms/s $\times 10^{8}$)
@@
- ZS, - 2D OM $1.6$ $1.3$ $19\,\%$
- ZS, ++ 2D OM $3.6$ $2.1$ $42\,\%$
++ ZS, - 2D OM $7.1$ 99 $27\,\%$
++ ZS, ++ 2D OM $15$ $13$ $13\,\%$
Table 1: A comparison of the 2D MOT loading rates obtained experimentally ($\mathcal{F}_{exp}$) and numerically ($\mathcal{F}_{sim}$). “++” and “-” refer to the 2D OM or the ZS magnets being used and not being used, respectively.
Starting with row/case 1, where the atom dynamics are not affected by the 2D OM or the ZS magnets, the discrepancy between the experiment and simulation is less than $20\,\%$. This is noteworthy considering there is only one free numerical parameter (the total atomic flux emitted by the oven). The simulation is highly influenced by the initial velocity profile, and by employing our small-angle approximation for the oven (Sec. II.A) proves to yield closely matching results. The atom lifetime is also an important factor to consider (see Appendix B), given that the atoms can get shelved to a metastable state as they travel through the Zeeman nipple and hence elude the 2D MOT capture.
@@
The 3D MOT field is realized with a pair of rectangular anti-Helmholtz coils surrounding the glass cell at the main science region [coils indicated by Fig. 1 but not shown in Fig. 2(a); $2\,\text{cm}\,\times\,2.5\,\text{cm}$ radial ($\textsf{x}\times\textsf{y}$) cell dimensions], resulting in an approximately cylindrically-symmetric field around the transfer axis (z axis) at the MOT center ($31.5$ cm below the 2D MOT center). Along this axis, we have two independent beams, each with a power of 3.5 mW and a $1/e^{2}$ radius of $3.2$ mm; one stems from below the 3D MOT with a fixed detuning of 46-46 MHz and the other counter propagates from above the 2D MOT with a detuning that we vary for testing our loading technique (Sec. III). We note that we also consider a special configuration where the bottom beam is slightly misaligned, as discussed later. In the radial plane (xy plane), there are two retroreflected beams with an individual power of 2 mW, a $1/e^{2}$ radius of $2.4$ mm, and a detuning of 30-30 MHz. They are lowered by approximately one beam radius with respect to the quadrupole field zero, resulting in increased loading efficiency (discussed in detail in Sec. III). Additionally, mirroring them at $25\,^{\circ}$ with respect to the radial axis perpendicular to the objectives axis (i.e., with respect to the y axis) grants the necessary optical access for the high NA optical tweezers while simultaneously covering most of the science region for enhanced loading (see the Supplemental Material [49] for the numerical verification). One of the retroreflected pairs also contains the repumping light with a total power of $8$ mW; we find the repumping to be critical as most of the atoms get shelved to the metastable state $5s5p\,^{3}\!P_{2}$ before reaching the science region.
The 3D MOT magnetic field is modeled by Eq. A6, which is an approximation to the field created by the rectangular coils used in our experiment. Note that the modeled field is present in the transfer region (the same is true for the 2D MOT field); this information feeds into the scattering cross sections (Eq. A11), thus affecting the atom dynamics. The radiation pressure forces corresponding to the radial plane beams take into account the lowering (1 radial beam radius) as well as the angling ($25\,^{\circ}$). The repumping is present in one of the radial beam pairs, like in the experiment. As mentioned in the introduction of Sec. II, when a given superparticle drifts into the repumping region, it is revived with a new lifetime that is enhanced compared to the non repumping region.
@@ III. Main results
Our 3D MOT loading involves a simple technique depicted  in  Fig.  1:  The  axial  3D MOT  laser  beams  are
used to create an atomic molasses beam from the location of the 2D MOT, providing cold atoms for the loading. This atomic beam is the result of a radiation-pressure imbalance between the axial beams, which is attained by (i) using unequal detunings in case they are counter propagating or (ii) having the bottom beam slightly angled to the 3D MOT when equally detuned. In the latter case, introducing a detuning difference results in enhanced loading (discussed in Sec. III.C).
With the moving molasses technique, the 3D MOT can be loaded for a broad range of parameters (discussed below), as the atoms accumulate below the quadrupole field zero as set by the axial pressure imbalance. This technique creates an advantageous MOT laser beam configuration as compared to adding a push or cold flux beam, which may otherwise require operating the MOT beams at angles compromising the loading [22, 23]. Moreover, a broad optical access to the scientific study region is achieved, creating additional laser-pathways for atom array manipulation with a high NA optical tweezer setup.
Below in this section, we first develop an analytical model for the 3D MOT loading using our technique and then compare the main experimental observations to this model as well as the numerical model. Lastly, we discuss the configuration with a misaligned bottom beam. We note that this special configuration obfuscates analytical treatment and is thus not central to our present analysis.
@@ A. The analytical model
Here we develop an analytical theory for the 3D MOT loading, with Fig. 1 providing a visual guide (the aligned case is treated). We perform a step-by-step analysis, starting with the 2D MOT region, then describe the molasses beam in the transfer region (between the 2D and 3D MOTs) and the atom capture in the 3D MOT, and finally obtain an equation for the atom number in the 3D MOT. The theory incorporates the atomic diffusion effect and the shelving to the metastable state $5s5p\,^{3}\!P_{2}$ during the transfer, as well as the loss to the dark state $5s5p\,^{3}\!P_{0}$ in the final trap region where repumping is present. Note that the lifetimes associated with these respective processes (shelving and loss) are intrinsically Poissonian (present in our numerical model), but here we choose average lifetimes for simplicity. Furthermore, we treat cut offs with Heaviside functions providing abrupt boundaries, which is a legitimate choice when considering dimensional constraints but crudely approximating transitions in a MOT that are naturally expected to be smooth. Achieving such smoothness through the complexification of our model is not necessary considering the close match with experiment (Sec. III.B).
Starting from the 2D MOT, the unequally detuned axial 3D MOT beams create an atomic molasses beam directed at the 3D MOT with the flow rate $\mathcal{F}_{0}$. In addition to having a strong longitudinal velocity component, the atoms in this beam naturally possess a transverse velocity distribution impacting their capture probability in the 3D MOT.
As the atoms continue downstream in the transfer region, they accumulate a radial position spread due to the diffusion stemming from the radiation pressure of the axial (transfer) beams. This spread can be affected by the shelving to the metastable state, which freezes the diffusive velocity distribution. To obtain the transfer probability $\mathcal{P}_{\oplus}$ into the 3D MOT due to the randomness in the atomic beam, we take the radial spread to be normally distributed with the variance $\sigma_{r}^{2}$ and integrate this distribution as follows. Two different regimes are considered with their corresponding integration limits, in order to accommodate for the substantial range of magnetic field gradients explored in the experiment. In the first regime, we assume that the gradient of the 3D MOT is shallow enough for its radial capture to extend beyond the cell bounds, and thus the integration limits are set from $d_{0}$ to $d_{0}$, where $d_{0}$ is the distance a radial beam travels inside the cell to its center (3D MOT). In the second regime, the situation is opposite, and thus the integration limits are from $d_{\textsf{c}}$ to $d_{\textsf{c}}$, where $d_{\textsf{c}}$ denotes the radial confinement distance. The integrations yield respective error functions erfer\!f for the two regimes and, by mathematically separating them with appropriate Heaviside functions $H$, yield
@@
$\hskip-85.0pt\begin{aligned} \mathcal{P}_{\oplus}&=H(d_{\textsf{c}}-d_{0})\times er\!f\left(\frac{d_{0}}{\sqrt{2\sigma_{r}^{2}}}\right)\\ &+H(d_{0}-d_{\textsf{c}})\times er\!f\left(\frac{d_{\textsf{c}}}{\sqrt{2\sigma_{r}^{2}}}\right)\end{aligned}$ (1)
where $H(0)=1/2$ for $d_{\textsf{c}}=d_{0}$ and continuity is preserved. The unknown variables are as follows:
@@
(i) $d_{\textsf{c}}=\frac{3|\delta_{0}|}{\mu B^{\prime}}$, where $\frac{\delta_{0}}{2\pi}$ is the radial beam detuning, $\mu$ is the gyromagnetic ratio, and $B^{\prime}>0$ is the axial gradient (twice stronger than the radial). This expression is obtained by assuming for simplicity that the furthest capture occurs midway between the peak capture distance $\frac{2|\delta_{0}|}{\mu B^{\prime}}$ and twice this distance (where the radiation pressure force is significantly diminished). We note that in general, the confinement distance has a complex dependence on different trapping parameters [60].
(ii) $\sigma_{r}=H(\tau-t)\sigma_{r,t}+H(t-\tau)\sigma_{r,s}$, where the Heaviside functions separate the cases where the time $\tau$ until the metastable-state shelving is longer than the transfer time $t$ and vice versa; and $\sigma^{2}_{r,t}$ and $\sigma^{2}_{r,s}$ are the variances of the Gaussians describing the radial position spreads for
@@
these respective cases. When the shelving occurs, the corresponding speed Gaussian is assumed to freeze at $\tau$ and evolve the position Gaussian until $t$ by being in quadrature with the position Gaussian at $\tau$ and through a correlation term; thus, $\sigma_{r,s}=\sqrt{\sigma^{2}_{r,\tau}+\sigma^{2}_{v,\tau}(t-\tau)^{2}+2\rho\sigma_{r,\tau}\sigma_{v,\tau}(t-\tau)}$, where $\sigma^{2}_{r,\tau}$ and $\sigma^{2}_{v,\tau}$ are the variances of respectively the position and speed Gaussians at $\tau$, and $\rho=\frac{\sqrt{3}}{2}$ is the correlation between them (Brownian in nature). The remaining quantities seen in $\sigma_{r}$ are expressed as follows: The transfer time $t=\frac{d-d_{\textsf{e}}}{-v_{max}}>0$, where $d$ is the distance between the 2D and 3D MOTs; $d_{\textsf{e}}=\frac{\delta_{\text{b}}-\delta_{\text{t}}}{2\mu B^{\prime}}$ is the axial equilibrium position of the 3D MOT, as we assume the bottom and top beams of respective detunings $\frac{\delta_{\text{b}}}{2\pi}$ and $\frac{\delta_{\text{t}}}{2\pi}$ have the same intensity; and $v_{max}=\frac{\delta_{\text{b}}-\delta_{\text{t}}}{2k}$ is the maximal speed achieved during the transfer, assuming there is no magnetic field in the transfer region ($k=\frac{2\pi}{\lambda}$ is the angular wavenumber of the blue transition). The time until shelving, $\tau$, is obtained using Eq. B1 with $\mathcal{E}=1$ (no repumping) and with $\Gamma_{23}^{\prime}$ as in Eq. B2, but here we use a two-level model total saturation parameter for the bottom and top beams, $s_{\text{bt}}=\frac{I_{\text{bt}}/I_{sat}}{1+4\frac{(\delta_{\text{b}}-kv_{max})^{2}}{\Gamma^{2}}}+\frac{I_{\text{bt}}/I_{sat}}{1+4\frac{(\delta_{\text{t}}+kv_{max})^{2}}{\Gamma^{2}}}$ ($I_{\text{bt}}$ is the intensity of an axial beam). The variance $\sigma^{2}_{r,\tau}=\frac{2D_{r}\tau^{3}}{3^{2}m^{2}}$ ($\sigma^{2}_{r,t}$ uses $\tau\rightarrow t$) and $\sigma^{2}_{v,\tau}=\frac{2D_{r}\tau}{3m^{2}}$, where the transfer-region diffusion coefficient $D_{r}$ is obtained using Eq. A13 with the total saturation parameter $s_{\text{bt}}$. Note, finally, that our definition of $\sigma_{r}$ assumes that the atomic beam originates from a single point. An improved assumption would be adding $\sigma_{r}$ in quadrature with the radial 2D MOT size, which is proportional to the square root of its temperature [61]. We omit this here, as our main results (Sec. III.B) are qualitatively unaffected by this change.
@@
Next, let us consider writing the probability $\mathcal{P}_{\ominus}$ incorporating various 3D MOT cut offs as well as the change in the confinement strength as the MOT equilibrium position $d_{\textsf{e}}$ shifts with respect to the fixed radial beams center. The first cut off we introduce ensures that the MOT is empty when $\delta_{\text{t}}$ is positive and larger than $|\delta_{\text{b}}|$, as the sign of the confinement is then flipped. Our next two cut offs ensure the MOT is empty when the radial beams do not encompass $d_{\textsf{e}}$ (as set by the axial beam imbalance), given the locations $d_{u}$ and $d_{l}$ of the upper and lower radial beam edges with respect to the quadrupole field zero. Note that additional cut offs can be introduced to ensure an empty MOT when the transfer speed overcomes the axial capture speed and when $d_{\textsf{e}}$ exceeds the axial confinement distance; we omit these here as for our parameters the first cut off condition is stronger. We represent the introduced cut offs (three in total) with  Heaviside functions.  For the change in the  confinement  strength  with
Figure 4: Diagrams showing the 3D MOT atom numbers (normalized; see text for absolute) using our loading technique connecting the 2D and 3D MOTs via moving molasses, obtained analytically, experimentally, and numerically. The experimental and numerical data points are obtained at the values seen on the axes, with the contours enhancing the features. The vertical dashed line marks the fixed bottom beam detuning (46-46 MHz), and the slanted dashed line delineates approximately the sloped analytical feature. The cross and the dotted circle highlight respectively the location with the greatest analytical atom number ($44.8$ G/cm gradient at 7.3-7.3 MHz top beam detuning) and the greatest experimental as well as numerical atom number ($48$ G/cm gradient at 10-10 MHz top beam detuning).
the shifting $d_{\textsf{e}}$, we assume a Gaussian behavior that depends on the trapped atom position $d_{p}$ with respect to the center of the radial beams ($w_{0}$ waist radius), where their intensity maximum is assumed to provide the strongest (unity) confinement. Thus,
$\hskip-95.0pt\hskip 33.0pt\begin{aligned} \mathcal{P}_{\ominus}&=H(-\delta_{\text{t}}-\delta_{\text{b}})H(d_{u}-d_{\textsf{e}})H(-d_{l}+d_{\textsf{e}})\\ &\times exp\left(-2\frac{d_{p}^{2}}{w_{0}^{2}}\right)\end{aligned}$ (2)
The unknown variables in this equation are as follows:
@@
For the final component in our analytical theory, we consider the time $\tau_{\text{loss}}$ until the dark-state decay occurs in the trap. We obtain it using Eq. B1 with $\mathcal{E}=27$ (repumping present) and with $\Gamma_{23}^{\prime}$ as in Eq. B2, but here we use a two-level model total saturation parameter for all the 3D MOT beams, $s_{\text{all}}=s_{\text{bt,B}}+4s_{0}$, where $s_{\text{bt,B}}$ is the total saturation parameter given by $s_{\text{bt}}$ with the substitution $kv_{max}\rightarrow\mu B^{\prime}d_{\textsf{e}}$ (assuming a motionless atom), and $s_{0}=\frac{\frac{I_{0}}{I_{sat}}exp\left(-2\frac{d_{p}^{2}}{w_{0}^{2}}\right)}{1+4\frac{\delta_{0}^{2}}{\Gamma^{2}}}$ is the radial beam saturation parameter, with $I_{0}$ the peak radial beam intensity.
Putting all of the above together, we obtain the following equation for the atom number in the 3D MOT (when the transfer from the 2D MOT location occurs with the flow rate $\mathcal{F}_{0}$):
$\hskip-100.0pt\begin{aligned} N=\frac{\mathcal{P}}{\Gamma_{\text{loss}}}\times\mathcal{F}_{0}\end{aligned}$ (3)
where $\mathcal{P}=\mathcal{P}_{\oplus}\times\mathcal{P}_{\ominus}$ combines the two probabilities $\mathcal{P}_{\oplus}$ and $\mathcal{P}_{\ominus}$   affecting   the overall   transfer   efficiency,   and
$\Gamma_{\text{loss}}=1/\tau_{\text{loss}}$ sets the dominant loss rate in the 3D MOT due to the dark-state decay. We note that $\Gamma_{\text{loss}}$ bounds the parameter space for optimal transfer, as discussed below.